# SoK: “Interoperability vs Security” Arguments: A Technical Framework

### v1

@@ 3.3 Hybrid Concerns
To open P2P Wi-Fi up more broadly, the Commission proposes [^22] DMA.100203 – Article 6(7) – Apple – iOS – SP – Features for connected physical devices that Apple either make AWDL available to third-parties or that Apple facilitate third-party interoperation through Wi-Fi Aware, [^23] Also called Neighbor Awareness Networking (NAN). which is a public protocol that has similar functionality and is based on AWDL[27]. If Apple makes AWDL public, there is little engineering to be done, as long as the protocol is indeed secure. If the security of AWDL is reliant upon its details remaining secret, i.e. security through obscurity, Apple and its users might be better served by Wi-Fi Aware or an improved version of AWDL.
@@ 3.3.1 Analysis: Hybrid Concerns
The hybrid cases are more than just the sum of the other two; we see patterns specific to this more complicated case.
1. How do interoperation and economic incentives clash (or not)? The platform’s market power is again a key feature of these cases. Since third-party products compete directly with those of the incumbent company (e.g., Apple Watch competing with third-party watches), the company has incentive to protect their current market share. At the same time, the company has control over access to those key functionalities and can block access for competitors. In contrast to vetting concerns, interoperation means that third parties gain access to functionality that was not previously available to third-party developers at all. Since the company must both build access to previously exclusive functionality and also vet who gets that access, the platform’s control over the situation is even greater than in engineering or vetting cases. Thus, the company’s ability to block access is more robust in the hybrid case.
2. How do security arguments and economic incentives align (or not)? The platform has a clear economic incentive to protect their exclusive products that are bundled or complementary to their main devices. Preventing third-party developers from offering the full functionality available on Apple’s own products reduces the value of those other products and puts theirs at an advantage, i.e., self-preferencing. Supporting interoperation would require security engineering in concert with third parties. These incentives coalesce into significant motivation for a platform to keep systems closed.
3. What needs to change in order for the company to support interoperation? Not only must the platform build secure APIs or other means of access for third-party developers, but many of these APIs may need to be bespoke for specific uses. This could be a significant engineering task. The platform can vet all proposals for interoperation; this enables them to enforce security standards, and keep tabs on what third-party developers are trying to do. Interoperation in hybrid cases is likely to be subject to more scrutiny from the platform in control than in vetting cases, and to require more coordination between the platform and third parties. Although perhaps costly, this may make for better, more secure products.
@@ 4 Comparative Analysis and Takeaways
We now offer a comparative analysis, identifying key considerations across our taxonomy and case studies. First, we consider economic incentives, market power, and security, corresponding to Analysis Questions 1 and 2. Then we discuss security trade-offs, corresponding to Analysis Questions 3 and 4. Finally, we offer big-picture takeaways. We also summarize our analysis in Table 2 in the Appendix.
@@
Market power is usually the purview of economists and lawyers, but recognizing the nature of a company’s market power can be a key step in understanding the ways in which security can become subordinated to other business goals.
In our framework, we do not categorize cases by the nature of the incumbent company’s market power, and yet different security concerns seem to correlate with different types of market power. For instance, we notice that cases with security engineering concerns are also those in which the market power of the incumbent is just its size. The relative size of a company, or, more precisely, its market share, and the resultant network effects, can confer significant power without more complicated ‘gatekeeper’ dynamics. As noted in Section 2, these instances are also often those in which the necessary interoperation is horizontal, suggesting that the parties in these cases might be on a more equal footing when it comes to making technical decisions.
The nature of the security concerns and the discourse around them is, in some sense, the simplest for security engineering concerns: something needs to be built. That does not mean that these concerns are always as insurmountable or substantiated as the companies are incentivized to suggest. There is an engineering task, but the technology to interoperate securely often exists or can be developed. It is also fairly clear what ‘interoperating securely’ will look like, in most cases, which stands in contrast to the other two categories.
This pattern has a counterpart in security vetting concerns, which involve instances of vertical interoperation (with a platform), as this type of argument is made by platforms vetting other products. As noted in Section 3, vetting concerns involve company discretion regarding security decisions and policy choices, and rarely have clear-cut engineering answers. Thus, vetting concerns have the potential to be more malleable, and therefore more easily tailored to fit the company’s economic needs. In addition to this malleability, the platform position gives the company outsized economic power in these scenarios, and these two aspects combine into significant power to impose punitive prices or exclude third-party developers. Thus, this type of concern should be identified as particularly susceptible to a blurring of boundaries between security and economic incentives, compared to engineering concerns.
The hybrid case is not simply a dilution of the dynamics of the other two cases. The platform power manifests here too. Rather than using platform power to extract fees, as with security vetting concerns, hybrid cases see companies using their platform power for self-preferencing their products, hobbling or sometimes entirely excluding potential competition. In both vetting and hybrid cases, the company acts as the platform gatekeeper. However, the extensive control of the incumbent company to block competition through both vetting and engineering concerns is distinct to hybrid cases, meaning the platform power is thus particularly entrenched.
@@
The engineering demands in a hybrid case are no less significant than in a security engineering case and the platform power is bolstered by legitimate engineering concerns. This means there are two lines of defense to keep out both bad actors and competitors, compounding the effects we see in pure engineering or vetting cases. Third-party competitors may be more completely shut out in this case as compared to the other two. Without interoperation, messaging apps can still function with their own users, and apps can be distributed on iPhones as long as they adhere to Apple’s guidelines and fee schemes. But without access to the HCE, there can be no alternative to Apple Pay on the iPhone. Third-party smartwatches can function without access to all hardware functionalities, but their utility may be significantly limited.
@@
Thus, hybrid security concerns should be evaluated with careful consideration of these unusually strong market ramifications as compared to engineering or vetting concerns.
Security trade-offs
@@
Security engineering involves many practical trade-offs: between security and efficiency, usability, profit, and more. For example, achieving usability and security at once is well known to be challenging [32]. Usability also interacts with companies’ economic incentives: a phone with such tight security that it is unusable would be hard to sell. Economic interests can also motivate pressures against interoperation. Just as undermining usability through a narrow focus on security can ruin a product, undermining competition through a narrow focus on security can ruin a market. A key difference is that the latter aligns with companies’ interests, so could conceivably happen, whereas the former is against their interests, so would never happen in practice.
The simplest case is again the security engineering concern. Inherent in the idea of an engineering concern is the idea that the problem can be solved by engineering. The trade-off, then, relates to how much to invest in solving the engineering problem and how much of a burden that is. In practice, of course, even the best, well-funded systems have bugs and vulnerabilities. Thus, companies must deploy systems that might be vulnerable. This tension carries over into interoperation, where again systems might have vulnerabilities despite their best efforts. As such, we see using security arguments to push against interpretation as often ultimately about business interests more than security. We distinguish the preceding type of statement from the seemingly rarer type of security engineering concern that states that “interoperation cannot be done securely at all as a technical matter,” which is a different matter that must be evaluated from a technical security perspective. [^24] The statement that E2EE messaging services cannot interoperate with unencrypted messaging services while maintaining their level of security would fall in this category. (Note: Regulators do not require such interoperation.) We believe it is essential to systematically distinguish between these, as market actors are incentivized to obfuscate the distinction between these two types of security engineering concerns and evoke the higher security stakes of the latter in the former.
@@
When companies raise security vetting concerns, they may argue that security considerations necessitate limitations on interoperation, or perhaps preclude interoperation entirely. Recall that security vetting concerns arise from company policies that do not permit interoperation or make it costly or difficult, rather than technical interoperability issues. Policies generally involve discretionary trade-offs between security and other interests. The trade-offs made in existing policies are an essential reference to evaluate how much of a divergence from existing security practices, and how much of an economic burden, a proposed policy change would be.
Consider the case of alternative app stores (Case Study 3.2). The status quo involves a profitable fee structure by a gatekeeper platform. The aim of the proposed interoperation is primarily to allow developers to escape the fees, not existing vetting guidelines. Much like the App Store Guidelines, both the guidelines for notarization and the potential guidelines of any alternative app store will consist of policy decisions. The interoperability mandate does not suggest Apple permit app stores with policies that are worse for security than current App Store Guidelines. Apple’s proposed solution involves preserving a fee structure profitable for Apple across all alternative app stores, the details of which would make operating on alternative app stores not worth it for many developers. One reason they give for the necessity of this fee structure is security. While security vetting can be important and does incur some costs, meaning some fee collection may be beneficial for security, the structure of this “security vs. interoperability” argument should highlight that it is not just about security, and raise questions around the (mis)alignment of economic incentives, security arguments, and interoperation.
@@
The hybrid case features all of the aforementioned types of trade-offs. That said, in some ways, the hybrid trade-offs are of a slightly different character than the engineering and vetting trade-offs. Those seeking to interoperate with OS functionality reserved for device manufacturer use are fewer than those seeking to make apps for smartphones, so vetting could be a project of a substantially smaller scale. Relatedly, interoperation could be bespoke, involve more monitoring, and require less scalability.
@@
Our framework, and our discussion of case studies through the lens of the framework, highlight how the following themes feature in “security vs. interoperability” discourse in impactful policy contexts in the wild.
@@
In instances in which vetting is the primary type of security, incumbent companies necessarily make trade-offs and also have strong platform control that can be leveraged to extract fees. The ‘hybrid’ combination of platform power and guarded access to functionality can grant incumbent companies even greater power to exclude their direct competition.
@@ 5 Conclusion
The security community has an important role to play in impactful ongoing competition policy debates due to the prevalence of security arguments deployed to argue against interoperation. Security arguments should always be taken seriously, considering both their security implications and their economic and policy implications together. In this paper, we have presented our framework for analyzing “security vs. interoperability” arguments within their broader economic and policy context, along with case studies and analysis illustrating the versatility, utility, and takeaways of our framework.
@@ Acknowledgments
The authors would like to extend special thanks to Mallory Knodel for helpful discussion and many insights. We are also grateful for the generous help of Alexander Heinrich, who provided indispensable knowledge on technical matters. Thanks also to Andrés Fábrega and Mathy Vanhoef.
@@
Table 2: Summary of analysis results

### v2

@@ 4.3 Hybrid Concerns
To open P2P Wi-Fi up more broadly, the Commission proposes [^35] DMA.100203 – Article 6(7) – Apple – iOS – SP – Features for connected physical devices that Apple either make AWDL available to third parties or that Apple facilitate third-party interoperation through Wi-Fi Aware, [^36] Also called Neighbor Awareness Networking (NAN). which is a public protocol that has similar functionality and is based on AWDL[27]. If Apple makes AWDL public, there is little engineering to be done, if the protocol is indeed secure. If the security of AWDL is reliant upon its details remaining secret, i.e. security through obscurity, Apple and its users might be better served by Wi-Fi Aware or an improved version of AWDL.
@@ 4.3.1 Analysis: Hybrid Concerns
The hybrid cases are more than just the sum of the other two; we see (arguably more complex) patterns specific to this category.
1. How do interoperation and economic incentives clash (or not)? The platform’s market power is again a key feature of these cases. Since third-party products compete directly with those of the incumbent company (e.g., Apple Watch competing with third-party watches), the company has incentive to protect their current market share. At the same time, the company has control over access to those key functionalities and can block access for competitors. In contrast to vetting concerns, interoperation means that third parties gain access to functionality that was not previously available to third-party developers at all. Since the company must both build access to previously exclusive functionality and also vet who gets that access, the platform’s control over the situation is broader than in either engineering or vetting cases. Thus, the company’s ability to block access may be more robust in the hybrid case.
2. How do security arguments and economic incentives align (or not)? The platform has a clear economic incentive to protect their exclusive products that are bundled with or complementary to their main devices. Preventing third-party developers from offering the full functionality available on Apple’s own products reduces the value of those other products and puts theirs at an advantage, i.e., self-preferencing. Supporting interoperation would require security engineering in concert with third parties. These incentives may coalesce into significant motivation for a platform to keep systems closed.
3. What needs to change in order for the company to support interoperation? Not only must the platform build secure APIs or other means of access for third-party developers, but many of these APIs may need to be bespoke for specific uses. This could be a significant engineering task. The platform can vet all proposals for interoperation; this enables them to enforce security standards, and keep tabs on what third-party developers are doing. Given the previously exclusive nature of the functionality opened up, interoperation in hybrid cases may be subject to more scrutiny from the platform in control than in vetting cases, and may require more coordination between the platform and third parties. Although perhaps costly, this may make for better, more secure products.
@@ 5 Comparative Analysis and Takeaways
We now offer a comparative analysis, identifying key considerations across our taxonomy and case studies. First, we consider economic incentives, market power, and security, corresponding to Analysis Questions 1 and 2. Then we discuss security trade-offs, corresponding to Analysis Questions 3 and 4. Finally, we offer big-picture takeaways.
Table III (Appendix) provides a succinct summary of our comparative analysis.
@@
Recognizing the nature of a company’s market power can be a key step in understanding the ways in which security can become subordinated to other business goals. There are two main types of market power that are relevant in the cases we discuss: regular market dominance, where a company is simply so large it dominates some market; and platform power, where incumbent companies are ‘gatekeepers’ of third parties’ access to customers. We did not specifically categorize cases by the nature of the incumbent company’s market power, and yet different security concerns seem to correlate with different types of market power. Cases with security engineering concerns are also those in which the market power of the incumbent is simply market dominance. The relative size of a company, or, more precisely, its market share, and the resultant network effects, can confer significant power without more complicated ‘gatekeeper’ dynamics. As noted in Section 3, these instances are also often those in which the necessary interoperation is horizontal, meaning third parties might be on a more equal footing when it comes to making technical decisions. Interoperating securely is generally an engineering task to be coordinated between the companies involved and, if there are existing standards, it is clear what it takes to ‘interoperate securely.’ We can see platform power clearly in security vetting concerns, which are also instances of vertical interoperation. As noted in Section 4, vetting concerns involve company discretion regarding security policy choices, and rarely have clear-cut engineering answers. As such, we view narratives around vetting concerns as potentially more easily tailored to fit a company’s economic needs. The platform position in these cases typically gives the company significant power to impose potentially punitive prices or exclude third-party developers. Thus, this type of concern should be identified as particularly susceptible to a blurring of boundaries between security and economic incentives, compared to engineering concerns. The hybrid case is not simply a dilution of the dynamics of the other two cases. Platform power manifests here too. Rather than using platform power to extract fees, as with security vetting concerns, hybrid cases more often feature or involve risks of companies using their platform power for self-preferencing their products, hobbling or entirely excluding potential competition. In both vetting and hybrid cases, the company acts as a gatekeeper. The extensive control of incumbent companies to block competition through both vetting and engineering is distinct to hybrid cases, meaning platform power is particularly entrenched.
@@
The engineering demands in a hybrid case appear in general to be no less significant than in a security engineering case. Thus, platform power is bolstered by legitimate engineering concerns. This means there are two lines of defense to keep out both bad actors and competitors, compounding the effects we see in pure engineering or vetting cases. Third-party competitors may be more completely shut out in hybrid cases as compared to the other two categories. Without interoperation, messaging apps can still function with their own users, and apps can be distributed on iPhones as long as they adhere to Apple’s guidelines and fee schemes. But without access to the HCE, there can be no alternative to Apple Pay on the iPhone. Third-party smartwatches can function without access to all hardware functionalities, but their utility may be significantly limited.
@@
Thus, hybrid security concerns should be evaluated with careful consideration of these multi-faceted market ramifications as compared to engineering or vetting concerns.
Security tradeoffs
@@
Security engineering involves many practical tradeoffs: between security and efficiency, usability, profit, and more. Achieving usability and security at once is well known to be challenging [31]. Usability also interacts with companies’ economic incentives: a phone with such tight security that it is unusable would be hard to sell. Economic interests can also motivate pressures against interoperation. Just as undermining usability through a narrow focus on security can ruin a product, undermining competition through a narrow focus on security can ruin a market. The simplest case is again the security engineering concern. Inherent in the idea of an engineering concern is the idea that the problem can be solved by engineering. The tradeoff, then, relates to how much to invest in solving the engineering problem and how much of a burden that is. In practice, of course, even the best, well-funded systems have bugs and vulnerabilities. Thus, companies must deploy systems that might be vulnerable. This tension carries over into interoperation, where, as in most other contexts, systems might have vulnerabilities despite their best efforts.
@@
When companies raise security vetting concerns, they may argue that security considerations necessitate limitations on interoperation, or perhaps preclude interoperation entirely. Recall that vetting is about interoperability policies, rather than technical interoperability. Policies generally involve discretionary tradeoffs between security and other interests. The tradeoffs made in existing policies are an essential reference to evaluate how much of a divergence from existing security practices, and how much of an economic burden, a proposed policy change would be.
@@
The hybrid case may feature all of the preceding types of tradeoffs. Yet the hybrid tradeoffs are of a slightly different character than the engineering and vetting trade-offs. Those seeking to interoperate with OS functionality reserved for device manufacturer use are fewer than those seeking to make apps for smartphones, so vetting could be a project of a substantially smaller scale. Relatedly, interoperation could be bespoke, involve more monitoring, and require less scalability.
@@
Our taxonomy, case studies, and analysis highlight how the following prominent themes feature in ‘security vs. interoperability’ discourse in impactful competition policy contexts in the wild.
@@
Where vetting is the primary type of security concern, incumbent companies necessarily make tradeoffs and also have strong platform control that can be leveraged to extract fees. The hybrid combination of platform power and guarded access to functionality can grant incumbent companies even more power to exclude their direct competition.
6 Future Directions and Policy Implications
Court proceedings and regulatory investigations in the area will continue to be common [64]. Many of the investigations we mention are ongoing (e.g., [35]). Our taxonomy and analytical framework are designed to help courts, policymakers, economists, and the security community better classify and analyze ‘security vs. interoperability’ arguments. Competition experts’ reasoning often focuses on key distinctions in economics and antitrust, such as between horizontal and vertical market integration. We draw connections between horizontal and vertical interoperation and engineering, vetting, and hybrid concerns, designed to support rigorous analysis of interoperation and related security issues in legal and regulatory settings. Our approach provides analytical tools to distinguish interoperability-related security concerns across diverse market patterns, and suggesting where SvI arguments need special regulatory and/or technical scrutiny, e,g., complex hybrid settings.
@@ 7 Conclusion
The security community has an important role to play in ongoing competition policy debates due to the prevalence of security arguments in apparent tension with interoperation. We have presented a systematization of “security vs. interoperability” arguments based on a systematic review of relevant EU cases, along with a taxonomy, analytical framework, and detailed case studies with comparative analysis.
@@ Acknowledgments
The authors would like to extend special thanks to Mallory Knodel for helpful discussion and many insights. We are also grateful for the generous help of Alexander Heinrich, who provided detailed consultation on certain technical issues. Thanks also to Michael A. Specter, Stephanie Chen, Andrés Fábrega, and Mathy Vanhoef.
@@
TABLE III: Summary of analysis results
Policy Area Legal Case Technology SvI Concern Type
DMA DMA.100097 Meta - Messenger E2EE Messaging Engineering
DMA DMA.100206 Apple - new business terms App Distribution and IAP Vetting
Alternative App Stores
Webkit
DMA DMA.100204 Apple - Physical Devices OS and Physcial Devices Hybrid
DMA DMA.100204 Apple - Process
Merger M.9660 Google / Fitbit
Antitrust AT.40452 Apple - Mobile payments NFC
Merger M.8306 Qualcomm / NXP Semiconductors
Antitrust AT.37792 Microsoft Windows OS No - Degrade
Merger M.8314 Broadcom / Brocade Fiber Channel SAN Switch
Merger M.3998 Axalto / Gemplus SIM chips No - Disclose
Merger M.5669 Cisco / Tandberg Telepresence Interoperability Protocol
Merger M.5984 Intel / McAfee Malware detection No - Both
Merger M.6564 ARM / Giesecke & Devrient / Gemalto / JV Trusted Execution Environment
Merger M.10806 Broadcom / VMware Hardware + Software for virtualization
TABLE IV: Methodology Results